OpenStudy (anonymous):

Which equations in point-slope form are equations of the line that passes through the point (4, 5) and (−3, −1)?

OpenStudy (anonymous):

@bohotness

OpenStudy (anonymous):

@bahrom7893

OpenStudy (anonymous):

@Mimi_x3

OpenStudy (bohotness):

yes

OpenStudy (anonymous):

can you help with this please

OpenStudy (bohotness):

sure

OpenStudy (anonymous):

okay thank you so much

OpenStudy (anonymous):

A: y - 3 7/6 (x - 1)

OpenStudy (anonymous):

B: y - 5 = 6/7 (x - 4)

OpenStudy (bohotness):

okay

OpenStudy (bohotness):

which one do you think ti is?

OpenStudy (anonymous):

C: y + 1 = 6/7 (x + 3)

OpenStudy (anonymous):

D: y - 4 = 7/6 (x - 5)

OpenStudy (anonymous):

idk i'm not sure

OpenStudy (anonymous):

at first i thought it was A tho

OpenStudy (dtan5457):

1.find slope 2.put in slope point formula

OpenStudy (anonymous):

but i'm not sure if i'm right

OpenStudy (dtan5457):

you know how to do that?

OpenStudy (bohotness):

okay -.- you need to find the slope if you dont known how we will show you how

OpenStudy (anonymous):

btw its choose all that are correct so there's more then one answer

OpenStudy (bohotness):

okay

OpenStudy (anonymous):

thank you :) i've been on this all niite

OpenStudy (dtan5457):

slope=\[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

yes i understand that

OpenStudy (dtan5457):

point slope formula: \[(y-y1)=m(x-x1)\]

OpenStudy (dtan5457):

plug the slope into the value of m, and you will get your answer

OpenStudy (anonymous):

k

OpenStudy (bohotness):

`~sighs ~ idtan got this so bye......

OpenStudy (anonymous):



OpenStudy (anonymous):

? thanx?

OpenStudy (dtan5457):

@bohotness you want to take it from here?

OpenStudy (dtan5457):

i'll hand you the medal of a life time if you do

OpenStudy (bohotness):

sure i guess but i am going to need some of your help :3

OpenStudy (anonymous):

umm -1-4 is 5 they have 6/

OpenStudy (dtan5457):

lol ok. go and let me see.

OpenStudy (bohotness):

thanks :3

OpenStudy (anonymous):

but they have 6/7 fractions -1-4 is 5 o.o

OpenStudy (anonymous):

not 6

OpenStudy (anonymous):

thats what i'm confused on

OpenStudy (anonymous):

and i guess i'm in this alone?

OpenStudy (anonymous):

thanx. anyway. you helpt plenty as i wanted? -.-

OpenStudy (bohotness):

no i am here go ahead and finsh posting